OpenStudy (anonymous):

54,55,59,61,61,62,68,70,72 how do i find the irst quartile, median, third quartile, and maximum of the data set?

OpenStudy (lynfran):

the first quartile can be found using this formula \[\frac{ 1 }{ 4 }(n+1)\]

OpenStudy (lynfran):

n=9

OpenStudy (anonymous):

9/4?

OpenStudy (lynfran):

since there 9 #s in the set

OpenStudy (anonymous):

Okay

OpenStudy (lynfran):

\[\frac{ 1 }{ 4 }(9+1)\]

OpenStudy (anonymous):

1/4(9+1

OpenStudy (anonymous):

1 can't be divided by 4 can it?

OpenStudy (anonymous):

Do i move the 9+1 over the 4.

OpenStudy (lynfran):

u will get \[\frac{ 1 }{ 4 }(10)\]\[2.5th value\] the average 2nd and 3rd values are 55 and 59 so \[\frac{ 55+59 }{ 2 }=?\]

OpenStudy (anonymous):

Lagging

OpenStudy (lynfran):

55+59/2=57 now 57 is ur 1st quartile value

OpenStudy (anonymous):

Okay. what is the 2.5th value?

OpenStudy (lynfran):

thats the average where u will find ur two numbers for the fisrt quartile

OpenStudy (lynfran):

so in this set of numbers the first quartile is between the the 2nd and 3rd #

OpenStudy (anonymous):

Okay

OpenStudy (lynfran):

so we add the 2nd and 3rd # and then divide the # by 2

OpenStudy (anonymous):

okay would that be 54 and 55?

OpenStudy (lynfran):

no thats the 1st and 2nd # ... u need the 2nd and 3rd #

OpenStudy (anonymous):

so 55 and 59?

OpenStudy (lynfran):

yes now add them

OpenStudy (anonymous):

100 +14 114

OpenStudy (lynfran):

59+55=114 now we divide by 2

OpenStudy (lynfran):

114/2=?

OpenStudy (anonymous):

57

OpenStudy (lynfran):

good now 57 is ur 1st quartile

OpenStudy (anonymous):

Alright.

OpenStudy (lynfran):

to find the median we located the middle value in the set....what # is in the middle?

OpenStudy (anonymous):

61?

OpenStudy (lynfran):

correct

OpenStudy (lynfran):

to find the 3rd quartile the formula is \[\frac{ 3 }{ 4 }(n+1)th . value \]

OpenStudy (lynfran):

\[\frac{ 3 }{ 4 }(9+1)th.value\]

OpenStudy (anonymous):

What do i do first add the 9+1 will give me 10 but what do i do next?

OpenStudy (lynfran):

\[\frac{ 3 }{4 }(10)th.value\]\[\frac{ 30 }{ 4 }th. value\]7.5 th. value so the 3rd quartile is between the 7th and 8th numbers

OpenStudy (anonymous):

Oh so we time it by 3 3*10 = 30 /4

OpenStudy (anonymous):

7th and 8th what does that mean?

OpenStudy (lynfran):

we locate the 7th and 8th #s in the data/set

OpenStudy (lynfran):

whats the 7th and 8 number frm the set

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

68 and 70

OpenStudy (lynfran):

good now add them and divide by 2

OpenStudy (anonymous):

138/2 = 69

OpenStudy (lynfran):

good thats ur 3rd quartile

OpenStudy (anonymous):

So we should be done.

OpenStudy (lynfran):

whats the maximum?

OpenStudy (anonymous):

Not sure. ^ ^;

OpenStudy (lynfran):

72..

OpenStudy (anonymous):

DX Ahh.

OpenStudy (anonymous):

Okay thanks.

OpenStudy (lynfran):

here more info on quartiles with a similar example http://www.mathsteacher.com.au/year9/ch17_statistics/06_quartiles/quartiles.htm

OpenStudy (lynfran):

welcome

OpenStudy (anonymous):

Thanks for the link too.

OpenStudy (lynfran):

the 2nd quatile \[\frac{ 2 }{ 4 }(10)th.value \]= 5th. value so the 5th value is 61 and yes 61 is the 2nd quartile

OpenStudy (anonymous):

Alright thanks.